OpenStudy (anonymous):

You have seen that the total number of lights in a triangular lighting rig is given by T(n) = n2/2 + n/2, while T(n+1) = n2/2 + 3n/2 + 1. If T(n) times T(n+1) = f(n), what is the reduced form of f(n)?

OpenStudy (anonymous):

any sketch posible

OpenStudy (anonymous):

Umm, let me get it

OpenStudy (anonymous):

Spotlight A lighting rig is a large bank of lights that creates lighting effects at concerts and sporting events. A company makes rigs with lights arranged in polygonal patterns. For example, a triangular rig may have 1, 3, 6, 10, or 15 lights, depending on its size. Thats all I have /:

OpenStudy (anonymous):

T(n)=n^2+n/2 Replace each n above with n+1 to obtain the value of T(n+1) T(n+1)=(n+1)^2+(n+1)/2 f(n)=T(n)*T(n+1)=(n^2+n/2)((n+1)^2+(n+1)/2) Expand the RHS above and simplify to obtain the reduced form of f(n)